# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

### v2

@@ Appendix C Hyper Parameters
In this section, we illustrate the detailed parameters adopted in this work, as shown in Table 4.
@@ Appendix D Data Construction
In this section, we introduce the construction details of TOMG-Bench and OpenMolIns dataset, as well as the prompt templates. Figure 4 demonstrates the overall data construction process for TOMG-Bench as well as the OpenMolIns. Generally, we utilize Zinc250K to construct the test samples in TOMG-Bench, while adopting PubChem for the training samples in OpenMolIns. Notably, when constructing the OpenMolIns, we exclude all the molecules in Zinc250K to avoid data leakage and ensure the novelty score of the generated molecules.
Figure 4: Data construction workflow of TOMG-Bench & OpenMolIns.
@@ D.2 MolOpt
MolOpt is designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task. Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time. In contrast, TOMG-Bench requires one single LLM to optimize molecules with different metrics and optimization directions. In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED. The prompt templates for MolOpt are listed in Table 7.
@@ Appendix E Testing Process
0:  Pass or Not (Bool) if $t$ is AddComponent then    if $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false else if $t$ is DelComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 then        return true     else        return false else if $t$ is SubComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 and $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false
@@
Algorithm 2 MolOpt Testing Process
0:  : Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Optimization Direction Requirement $R$
0:  : Pass or Not (Bool) if $t$ is LogP then    if $LogP$($m^{g}$) $>$ $LogP$($m^{o})$ and R is $\uparrow$ then        return true     else if $LogP$($m^{g}$) $<$ $LogP$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is MR then    if $M\!R$($m^{g}$) $>$ $M\!R$($m^{o})$ and R is $\uparrow$ then        return true     else if $M\!R$($m^{g}$) $<$ $M\!R$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is QED then    if $Q\!E\!D$($m^{g}$) $>$ $Q\!E\!D$($m^{o})$ and R is $\uparrow$ then        return true     else if $Q\!E\!D$($m^{g}$) $<$ $Q\!E\!D$($m^{o})$ and R is $\downarrow$ then        return true     else        return false
@@
0:  : Generated Molecule $m^{g}$, Subtask $t$, Atom List $A$, Bond List $B$, Functional Group List $G$, Requirements $R$
0:  : Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for $atom$ in $A$ do         if $|atom|$($m^{g}$) $\neq$ $R$[$atom$then            flag = false else if $t$ is BondNum then    for $bond$ in $B$ do         if $|bond|$($m^{g}$) $\neq$ $R$[$bond$then            flag = false else if $t$ is FunctionalGroup then    for $group$ in $G$ do         if $|group|$($m^{g}$) $\neq$ $R$[$group$then            flag = false return flag
@@ Appendix F Detailed Results
In this section, we first show the leaderboard of TOMG-Bench in Table 12, where Claude-3.5 achieves first place with an average weighted success rate of 35.92%. Notably, via instruction tuning on our OpenMolIns dataset, Llama-3.1-8B achieves 6th place, which outperforms all the existing open-source general LLMs.
Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
GPT-4o (OpenAI, 2023) - 49.08 32.29
Claude-3 (Anthropic, 2024a) - 46.14 30.47
Llama-3.1-8B (OpenMolIns-large) 8 43.1 27.22
Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
Llama-3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
Table 12: Leaderboard of TOMG-Benchmark.
Then, we compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5. Obviously, it can be concluded that Claude-3.5 outperforms almost all the other LLMs in these subtasks, except for the DelComponent task and the BondNum task. In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5, while Llama-3.1-8B (OpenMolIns-large) achieves the best performance in the BondNum task.
@@
Moreover, in the MolOpt task, QED, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization. In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
Figure 5: The radar maps compare the performance ($\overline{W\!S\!R}$) of some representative LLMs from different model categories in each subtask of TOMG-Bench.
Finally, we present the detailed experimental results of all the subtasks in Table 13, 14, and 15.
@@
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0374 0.5343 0.0200 0.1982 0.0768 0.5750 0.0442 0.3028 0.0102 0.3671 0.0037 0.1468
@@
Llama-3.2-1B (OpenMolIns-large) 0.1756 0.5676 0.0997 0.3216 0.1816 0.4963 0.0901 0.2466 0.0844 0.5415 0.0457 0.2958
Llama-3.1-8B (OpenMolIns-large) 0.5822 0.6541 0.3808 0.6730 0.5104 0.5074 0.2590 0.6896 0.5440 0.6258 0.3404 0.8400
@@
Table 13: Detailed results on MolEdit. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
@@
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0644 0.5055 0.0326 0.1664 0.0822 0.4410 0.0363 0.1604 0.0714 0.4757 0.0340 0.1796
@@
Llama-3.2-1B (OpenMolIns-large) 0.2898 0.5951 0.1725 0.3850 0.2644 0.5956 0.1575 0.3678 0.1996 0.5849 0.1167 0.3490
Llama-3.1-8B (OpenMolIns-large) 0.8054 0.6678 0.5378 0.8720 0.7122 0.6548 0.4663 0.8514 0.5224 0.6398 0.3342 0.8802
@@
Table 14: Detailed results on MolOpt. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
@@
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0040 0.6807 0.0027 0.1850 0.0080 0.7465 0.0060 0.2226 0.0008 0.7461 0.0006 0.2818
@@
Llama-3.2-1B (OpenMolIns-large) 0.0144 0.6490 0.0093 0.5616 0.0350 0.6150 0.0215 0.6186 0.0252 0.6373 0.0161 0.4412
Llama-3.1-8B (OpenMolIns-large) 0.0136 0.6634 0.0090 0.7582 0.0544 0.6614 0.0360 0.7456 0.1344 0.6396 0.0860 0.6435
@@
Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).

### v3

@@ Appendix C Hyper Parameters
[…]
@@
epochs(light) 10
epochs(small, medium) 5
epochs(large, xlarge) 3
@@ Appendix D Data Construction
In this section, we introduce the construction details of S2-Bench and OpenMolIns dataset, as well as the prompt templates. Figure 4 demonstrates the overall data construction process for S2-Bench as well as the OpenMolIns. Generally, we utilize Zinc250K to construct the test samples in S2-Bench, while adopting PubChem for the training samples in OpenMolIns. Notably, when constructing the OpenMolIns, we exclude all the molecules in Zinc250K to avoid data leakage and ensure the novelty score of the generated molecules.
Figure 4: Data construction workflow of S2-Bench & OpenMolIns.
@@ D.2 MolOpt
MolOpt is designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task. Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time. In contrast, S2-Bench requires one single LLM to optimize molecules with different metrics and optimization directions. In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED. The prompt templates for MolOpt are listed in Table 7.
@@
Among the three subtasks, the QED optimization, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization. In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
@@ Appendix E Testing Process
0: Pass or Not (Bool) if $t$ is AddComponent then    if $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false else if $t$ is DelComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ - 1 then        return true     else        return false else if $t$ is SubComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ - 1 and $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false
Algorithm 2 MolOpt Testing Process
0: Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Optimization Direction Requirement $R$
0: Pass or Not (Bool) if $t$ is LogP then    if LogPLogP($m^{g}$) >> LogPLogP($m^{o})$ and R is $\uparrow$ then        return true     else if LogPLogP($m^{g}$) << LogPLogP($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is MR then    if MRM\!R($m^{g}$) >> MRM\!R($m^{o})$ and R is $\uparrow$ then        return true     else if MRM\!R($m^{g}$) << MRM\!R($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is QED then    if QEDQ\!E\!D($m^{g}$) >> QEDQ\!E\!D($m^{o})$ and R is $\uparrow$ then        return true     else if QEDQ\!E\!D($m^{g}$) << QEDQ\!E\!D($m^{o})$ and R is $\downarrow$ then        return true     else        return false
@@
0: Generated Molecule $m^{g}$, Subtask $t$, Atom List AA, Bond List $B$, Functional Group List $G$, Requirements $R$
0: Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for atomatom in AA do         if |atom||atom|($m^{g}$) $\neq$ RR[atomatomthen            flag = false else if $t$ is BondNum then    for bondbond in $B$ do         if |bond||bond|($m^{g}$) $\neq$ RR[bondbondthen            flag = false else if $t$ is FunctionalGroup then    for groupgroup in $G$ do         if |group||group|($m^{g}$) $\neq$ RR[groupgroupthen            flag = false return flag
@@ Appendix F Detailed Results
In this section, we first show the leaderboard of S2-Bench in Table 12, where Claude-3.5 achieves the second place with an average weighted success rate of 35.92%. Notably, via instruction tuning on our OpenMolIns dataset, Llama3.1-8B achieves the first place, which outperforms all the other LLMs.
Rank Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
1 Llama3.1-8B (OpenMolIns-xlarge) 8 58.79 39.33
2 Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
3 Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
4 GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
5 GPT-4o (OpenAI, 2023) - 49.08 32.29
6 Claude-3 (Anthropic, 2024a) - 46.14 30.47
7 Llama3.1-8B (OpenMolIns-large) 8 43.1 27.22
8 Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
9 Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
10 Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
11 Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
12 GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
13 Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
14 Gemma3-12B (Team et al., 2025) 12 26.28 15.00
15 Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 7 25.07 14.61
16 Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
17 Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
18 chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
19 Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
20 Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
21 yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
22 Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
23 BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
24 MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
25 Llama3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
26 MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
27 MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
28 Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
Table 12: Leaderboard of S2-Bench.
Figure 5: Performance Comparison across several representative LLMs on S2-Bench.
We also compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5. Among the proprietary models, it can be concluded that Claude-3.5 achieves a powerful performance in these subtasks, except for the DelComponent task and the BondNum task. In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5. More importantly, we can also observe that Llama3.1-8B (OpenMolIns-xlarge) outperforms most of the other LLMs across these nine subtasks, achieving the SOTA performance.
@@
Finally, we also include a reasoning LLM, Deepseek-R1-distill-Qwen-7B (Guo et al., 2025), as shown in Table 12, which achieves the best performance among 7 and 8B level open-source general LLMs, showing the potential of LLM reasoning in open-domain natural language-driven molecule generation tasks.
Finally, we present the detailed experimental results of all the subtasks in Table 13, 14, and 15. Note that we do not provide results of vanilla galactica-125M, for it’s not instruction tuned and cannot respond to human instructions. All results of vanilla galactica-125M should be close to zeros.
@@
Llama3.2-1B-Instruct (Dubey et al., 2024) 0.0374 0.5343 0.0200 0.1982 0.0768 0.5750 0.0442 0.3028 0.0102 0.3671 0.0037 0.1468
Gemma3-12B (Team et al., 2025) 0.4300 0.7081 0.3045 0.6324 0.4948 0.5158 0.2552 0.7168 0.4208 0.7154 0.3010 0.7616
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.2667 0.6904 0.1841 0.0090 0.6367 0.5921 0.3770 0.0736 0.2156 0.7380 0.1591 0.0356
@@
Llama3.2-1B (OpenMolIns-large) 0.1756 0.5676 0.0997 0.3216 0.1816 0.4963 0.0901 0.2466 0.0844 0.5415 0.0457 0.2958
Llama3.1-8B (OpenMolIns-large) 0.5822 0.6541 0.3808 0.6730 0.5104 0.5074 0.2590 0.6896 0.5440 0.6258 0.3404 0.8400
Llama3.1-8B (OpenMolIns-xlarge) 0.7790 0.6769 0.5273 0.9468 0.8640 0.6166 0.5327 0.9132 0.6100 0.7434 0.4535 0.9596
@@
Table 13: Detailed results on MolEdit. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
@@
Llama3.2-1B-Instruct (Dubey et al., 2024) 0.0644 0.5055 0.0326 0.1664 0.0822 0.4410 0.0363 0.1604 0.0714 0.4757 0.0340 0.1796
Gemma3-12B (Team et al., 2025) 0.4528 0.694 0.3142 0.5912 0.3256 0.3548 0.1155 0.4478 0.1666 0.0869 0.0145 0.3322
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.4048 0.4416 0.1788 0.5622 0.2896 0.4535 0.1313 0.2034 0.0835 0.4556 0.0380 0.6074
@@
Llama3.2-1B (OpenMolIns-large) 0.2898 0.5951 0.1725 0.3850 0.2644 0.5956 0.1575 0.3678 0.1996 0.5849 0.1167 0.3490
Llama3.1-8B (OpenMolIns-large) 0.8054 0.6678 0.5378 0.8720 0.7122 0.6548 0.4663 0.8514 0.5224 0.6398 0.3342 0.8802
Llama3.1-8B (OpenMolIns-xlarge) 0.8822 0.6662 0.5877 0.9314 0.6982 0.6693 0.4673 0.9422 0.8648 0.6736 0.5825 0.9310
@@
Table 14: Detailed results on MolOpt. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
@@
Llama3.2-1B-Instruct (Dubey et al., 2024) 0.0040 0.6807 0.0027 0.1850 0.0080 0.7465 0.0060 0.2226 0.0008 0.7461 0.0006 0.2818
Gemma3-12B (Team et al., 2025) 0.0460 0.6017 0.0277 0.3414 0.0280 0.5973 0.0167 0.3318 0.0006 0.5940 0.0004 0.3224
Deepseek-R1-distill-Qwen-7B (Guo et al., 2025) 0.1574 0.6611 0.1041 0.4815 0.1631 0.7264 0.1185 0.5758 0.0385 0.6248 0.0241 0.5634
@@
Llama3.2-1B (OpenMolIns-large) 0.0144 0.6490 0.0093 0.5616 0.0350 0.6150 0.0215 0.6186 0.0252 0.6373 0.0161 0.4412
Llama3.1-8B (OpenMolIns-large) 0.0136 0.6634 0.0090 0.7582 0.0544 0.6614 0.0360 0.7456 0.1344 0.6396 0.0860 0.6435
Llama3.1-8B (OpenMolIns-xlarge) 0.1186 0.6834 0.0811 0.8540 0.1196 0.6746 0.0807 0.9000 0.3548 0.6393 0.2268 0.9492
@@
Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight the best and the second-best success rate (SR), as well as the weighted success rate (WSR).
Appendix G Limitations
Although S2-Bench is carefully designed and well-validated through our experiments, we still observe several limitations:
Data Distribution. In our data construction process, we allocate distributions to atoms, bonds, and functional groups with the aim of making our benchmark more reflective of real-world distributions. Nevertheless, the distribution we use is largely empirical and may not be sufficiently accurate to reconstruct real-world scenarios accurately.
The Selection of Subtasks. For the selection of subtasks, there are various choices in MolOpt, we compeletly understand that there are multiple viable options, especially when there are enormous metrics for evaluating molecular properties. However, we need to 1) consider the consistency between the three major tasks in S2-Bench and balance their weights (which means we can not select too many metrics for MolOpt, otherwise it would make the benchmark solely an optimization benchmark), and 2) ensure that the metrics can be efficiently and robustly calculated. Therefore, we only select the three common metrics, LogP, MR, and QRD, for MolOpt.
Appendix H Case Study
Finally, we include the case study to better illustrate the performance gap among different models.
Figure 6: Case study of AddComponent. In this example, we require LLMs to add a hydroxyl to the given molecule. Here, GPT-4o and Claude-3.5 barely modified the molecule and failed to add a hydroxyl to the given molecule, while other four models successfully passed the testing process. However, it can be clearly noted that although BioT5-base passed the testing process, the generated molecule of BioT5-base could be quite different from the original molecule.
Figure 7: Case study of DelComponent. In this example, we require LLMs to remove an amine from the given molecule. Here, GPT-4o, Claude-3.5, and BioT5-base all removed two amines, while the remaining models successfully passed the testing process.
Figure 8: Case study of SubComponent. In this example, we require LLMs to substitute the hydroxyl from the given molecule by a aldehyde. Here, although BioT5-base added a aldehyde, it failed to remove the hydroxyl. In contrast, the remaining models successfully passed the testing process.
Figure 9: Case study of LogP. In this example, we require LLMs to modfify the molecule to have a higher LogP value. Here, after the calculation of LogP values, we found all the generated molecules satisfy the requirement.
Figure 10: Case study of MR. In this example, we require LLMs to modify the molecule to have a higher MR value. Here, Llama3.1-8B-Instruct and Llama3.1-8B-Instruct (OpenMolIns-large) surprisingly failed the testing process. From the two errors, we could observe that the two models were trying to introduce extra rings to increase the MR value, while the ring numbers were misassigned.
Figure 11: Case study of QED. In this example, we require LLMs to lower the QED value of the given molecule. Compared to LogP and MR, QED is more difficult to predict. We could observe that GPT-4o, BioT5-base, Llama3.1-8B-Instruct, and Llama3.1-8B-Instruct (OpenMolIns-large) all increased the QED value of the molecule, while Claude-3.5 and Galactica-125M (OpenMolIns-xlarge) successfully captured the requirement and lowered the QED value.
Figure 12: Case study of AtomNum. In this example, we require LLMs to generate a molecule with 15 carbon atoms and 2 bromine atoms. Here, GPT-4o, Claude-3.5, BioT5-base, and Llama3.1-8B-Instruct all generated incorrect numbers of the specified atoms.
Figure 13: Case study of BondNum. In this example, we require LLMs to generate a molecule with 6 aromatic bonds, which normally exist in benzene rings. Here, GPT-4o, Claude-3.5, Llama3.1-8B-Instruct (OpenMolIns-large), and Galactica-125M (OpenMolIns-xlarge) captured the concept and generated molecules with a benzene ring in their structures. In contrast, BioT5-base failed to identify the requirement and Llama3.1-8B-Instruct generated too may aromatic rings.
Figure 14: Case study of FunctionalGroup. In this example, we require LLMs to generate a molecule with 2 benzene rings. Here, only Claude-3.5 failed to generate the correct number of benzene rings.